OpenStudy (mrs.ambrose614):

@elsa213 @sweetburger @will.h @glitter @awolflover @calculusxy CAN YOU GUYS HELP ME WITH THIS PROBLEM PLEASE. The graph below plots a function f(x): graph of line segment going through ordered pairs 0,150 and 3, 0 If x represents time, the average rate of change of the function f(x) in the first three seconds is ___.

OpenStudy (will.h):

considering the following coordinates (0,150) (3,0) Finding the rate of change of the function in the 1st 3 seconds use the rate of change law rate of change = \[\frac{ y2-y1 }{ x2-x1 }\] (it is also reprent the slope if this is a linear function) rate of change = \[\frac{ 0 - 150 }{ 3 - 0 }\] rate of change = \[-50\] so we just found the slope all left is to find the y intercept we could use y - y1 = m(x - x1) where m is the slope substitute y - 150 = -50(x - 0) y - 150 = -50x+ 0 y = -50x + 150 Thus the graph would look like this

OpenStudy (mrs.ambrose614):

Thank you will so the awnser would be -50????

OpenStudy (will.h):

for the rate of change yes but i thought you also needed to graph it.. so i made my way to get to the whole equation

OpenStudy (will.h):

oh i hadn't notice... there should have been a graph for you to post.. you didn't... you just needed the rate of change.. yes it would be -50

OpenStudy (mrs.ambrose614):

Thanks so much mr.will.h I hope you wouldn't mind helping me out with a few more sir??? @Will.H

OpenStudy (will.h):

not at all post them up in a new thread and tag me you welcome